bill invested $30,000 in two seperate accounts paying 3% and 4.1% annual simple interest. if bill earned a total of of $1091.73 from the two investments, how much was invested in each account.
I have no clue how to come up with the equation of this problem. The instructions say to find the equation for this investments problem andsolve it.

Answer by srinivas123(23)

let one amount be x then the second amount is 30000-x
Interest on first ammount i=ptr/100
(x*1*3 )/ 100. ----------------------equation (1)
Interest on second amount i=((30000-x)*1*4.1 )/ 100-------------equaion (2)
so total interest is equation (1) + (2)
3x/100 + 4.1(30000-x)/100 = 1091.73
after cross multiplication we get
3x+4.1(30000-x)= 109173
3x+123000-4.1x = 109173
-1.1x = -13827
------> x=12570
hence first amount is $12570 and second amount is $17430.00.